# Go Beyond Earth: Understanding Human Actions and Scenes in Microgravity Environments

### v3

@@ Abstract
Despite substantial progress in video understanding, most existing datasets are limited to Earth’s gravitational conditions. However, microgravity alters human motion, interactions, and visual semantics, revealing a critical gap for real-world vision systems. This presents a challenge for domain-robust video understanding in safety-critical space applications. To address this, we introduce MicroG-4M, the first benchmark for spatio-temporal and semantic understanding of human activities in microgravity. Constructed from real-world space missions and cinematic simulations, the dataset includes 4,759 clips covering 50 actions, 1,238 context-rich captions, and over 7,000 question–answer pairs on astronaut activities and scene understanding. MicroG-4M aims to support three core tasks: fine-grained multi-label action recognition, temporal video captioning, and visual question answering, thereby enabling a comprehensive evaluation of both spatial localization and semantic reasoning in microgravity contexts. We establish baselines using state-of-the-art models. All data, annotations, and code are available at https://github.com/LEI-QI-233/HAR-in-Space.
@@ 1 Introduction
Yuri Gagarin’s historic flight in 1961 marked the beginning of human space exploration. Since then, significant milestones have been achieved, including crewed lunar landings, the continuous operation of the International Space Station (ISS) for over $25$ years, and the participation of more than $650$ individuals in space missions Moskowitz & Wolf (2025). With numerous planned crewed missions in the near future, the frequency and complexity of human activities in space are expected to increase substantially. In this context, ensuring the safety, enhancing the operational efficiency of space missions, and safeguarding the health and well-being of astronauts are of paramount importance.
With the rapid advancement of artificial intelligence, the integration of robotic systems aboard spacecraft is anticipated in the near future. These systems will assist astronauts in routine and mission-critical tasks. Consequently, there is a growing demand for the development of deep learning-based scene understanding and action recognition methods tailored specifically for the unique challenges posed by the microgravity environment.
Human action recognition Wang et al. (2023a); Feichtenhofer (2020); Feichtenhofer et al. (2019), scenario captioning Chen et al. (2024), and Visual Question Answering (VQA) Hu et al. (2023) are essential for intelligent human-robot collaboration, particularly in space, where precise perception and understanding of astronauts’ actions and their surrounding context are crucial for ensuring operational safety, efficiency, and autonomous assistance under constrained conditions. This capability is crucial for ensuring mission efficiency, enhancing astronaut safety, and providing autonomous assistance in the confined and complex conditions of space habitats. Numerous human action recognition datasets and video caption datasets have been developed, playing an important role in various research areas within computer vision and in industrial applications Camarena et al. (2023); Pareek & Thakkar (2021); Al-Faris et al. (2020); Le et al. (2022).
However, most of the existing datasets for video scenario captioning and action recognition are recorded on Earth without microgravity settings, e.g., Kinetics400 Kay et al. (2017), AVA Gu et al. (2018), FineGym Shao et al. (2020), and ActivityNet Captions Chen et al. (2019).
@@
To address this gap, we introduce MicroG-4M, a new video benchmark specifically designed for spatio-temporal and semantic understanding of human activities in microgravity. The name “4M” reflects four characteristics: Multi-source (Real mission footage and physically plausible film), Multimodal (RGB + text annotations), Multi-task (HAR, captioning, VQA), and Microgravity.
@@
MicroG-4M comprises $4,759$ three-second video clips drawn from public YouTube footage of real space missions and carefully selected, realistic space-themed films to augment scenario diversity and coverage. The clips span scenes inside the ISS, the Tiangong Space Station, crewed spacecraft cabins, and extravehicular activities. The corpus contains more than $390,000$ annotated frames with bounding boxes for each visible individual and $13,000+$ action labels covering $50$ distinct actions. In addition, it includes $1,238$ descriptive clip captions created by human annotators and $7,000+$ open-ended question–answer pairs aimed at assessing factual, causal, and counterfactual understanding of scenes in microgravity. Both captions and QA annotations were carefully curated through multiple refinement rounds and data selection stages, ensuring high semantic fidelity, contextual accuracy, and relevance to the microgravity setting. Together, these resources enable multi-label spatio-temporal detection, fine-grained action recognition, caption generation, and VQA within a single dataset. To evaluate existing methods, we build the first comprehensive benchmark for these tasks, dubbed MicroG-Bench. We evaluate representative video encoders (SlowFast Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), and MViT Li et al. (2022)) for human action recognition and leading vision-language models (e.g., InternVideo Wang et al. (2022), Gemini 1.5 Pro Anil et al. (2023), GPT-4o Hurst et al. (2024)) for captioning and VQA tasks. Across all evaluated subtasks, state-of-the-art terrestrial models experience significant performance degradation, underscoring the unique challenges posed by microgravity environments, such as arbitrary orientations, floating objects, and confined spacecraft interiors. These results highlight the necessity of dedicated benchmarks to facilitate the advancement of more robust and generalizable AI systems tailored for space applications. MicroG-Bench, therefore, provides a unified, fair yardstick that will guide the development of more robust and generalizable perception and language systems for astronaut assistance and autonomous mission operations. We summarize our contributions as follows. We present MicroG-4M, the first video-based dataset for activity recognition and scene understanding specifically designed for the microgravity environment. Additionally, we establish the first benchmark to evaluate state-of-the-art human action recognition methods and vision-language models for fine-grained action recognition and video captioning/QA in this unique setting.
@@ 2 Related Work
Datasets for Action Detection, Video Captioning, and VQA. Progress in video understanding has been largely enabled by the introduction of benchmark datasets across action detection, captioning, and VQA. For action detection, early datasets such as UCF101 Soomro et al. (2012) and HMDB51 Kuehne et al. (2011) provided trimmed classification tasks, later extended by Kinetics Kay et al. (2017) and ActivityNet Caba Heilbron et al. (2015) to large-scale, untrimmed settings with diverse action categories. AVA Gu et al. (2018) further introduced spatio-temporal annotations for atomic actions, enabling fine-grained multi-label detection in realistic scenes. Video captioning evolved from MSVD Chen & Dolan (2011) and MSR-VTT Xu et al. (2016), which paired short clips with multiple sentences, to dense captioning in untrimmed videos via ActivityNet Captions Krishna et al. (2017) and procedural datasets like YouCook2 Zhou et al. (2018). Multilingual benchmarks such as VATEX Wang et al. (2019) expanded the task to cross-lingual settings with high-quality parallel annotations. In VQA, static image datasets like VQA v2.0 Goyal et al. (2017) and CLEVR Johnson et al. (2017) laid the foundation for compositional reasoning, while TGIF-QA Jang et al. (2017), MovieQA Tapaswi et al. (2016), and TVQA Lei et al. (2018) introduced temporal and multimodal reasoning in video. Recent efforts such as NExT-QA Xiao et al. (2021) and CLEVRER Yi et al. (2020) focus on causal inference and counterfactual reasoning. However, existing datasets are constrained to terrestrial environments and lack domain-specific complexities inherent to space-based activities. To bridge this gap, we introduce the first benchmark for video captioning and VQA in microgravity, enabling the evaluation of vision-language models in long-horizon, space-relevant scenarios.
@@ 3 Collection Methodology
We aim to construct a dataset comprising authentic footage recorded aboard the International Space Station and other spacecraft, as well as films that realistically depict microgravity conditions. In the following, we introduce our comprehensive pipeline for collecting, assembling, filtering, and annotating the MicroG-4M dataset, which supports multi-label spatio-temporal action detection, video captioning, and video question answering tasks from Internet-sourced videos.
Raw Video Information Collection. Video sources primarily include genuine microgravity footage from actual spacecraft missions and selected cinematic clips known for their realistic depictions of weightlessness. Authentic spacecraft videos were mainly retrieved from online video platforms, while cinematic clips were manually chosen based on their fidelity to real microgravity conditions. All videos were standardized to a resolution of $480p$ to maintain consistency. The final dataset comprises approximately $5,000$ three-second clips, predominantly composed of authentic spacecraft footage.
Dataset Assembly Pipeline. We established an automated pipeline to preprocess and structure raw videos, enabling consistent and accurate downstream annotation. The pipeline consists of three main stages: video segmentation, filtering, and automated bounding-box annotation.
@@
Filtering. Video clips undergo automatic filtering based on person detection and scene-transition analysis. Specifically, we employed YOLOv11 Khanam & Hussain (2024) for human detection and PySceneDetect Castellano to identify abrupt scene changes, discarding clips with insufficient human-action content or disrupted temporal continuity.
Bounding-box Annotation. Person bounding boxes are automatically annotated using YOLOv11 Khanam & Hussain (2024) detection combined with BoT-SORT tracking Aharon et al. (2022). To enhance annotation accuracy, adaptive strategies were employed by assessing video motion intensity using sparse optical flow methods Jeannin & Divakaran (2001); Ali (2013); Szeliski (2010). Annotation parameters were dynamically adjusted accordingly to optimize computational efficiency and annotation precision. The resulting structured annotations include bounding-box coordinates, unique identities, and detection confidence scores.
@@
Caption and VQA Annotation. For caption annotation, we developed a rigorous protocol that combines detailed visual analysis of video content with supplementary contextual information sourced from official aerospace agency documentation to ensure both accuracy and authority. Specifically, we compiled and referenced mission objectives, crew rosters, and official reports from various space agencies across different countries and missions, enabling precise identification of individual astronauts. Furthermore, we utilized historical spacecraft layout diagrams and functional distribution maps of the spacecraft cabins to accurately determine the astronaut’s location within the spacecraft as well as the equipment inside it. In parallel, annotators conducted thorough reviews of the complete video sequences, identifying subtle differences between adjacent frames to precisely describe astronauts’ actions, physical appearances, and interactions with the environment. These high-quality captions provide strong semantic grounding to support downstream tasks such as action recognition and context-aware retrieval.
For Video Question Answering (VQA), we adopted a structured approach based on Heilman & Smith (2009), formulating diverse questions using standard interrogative forms. Based on the complexity of the video captions, Large Language Models (LLMs) generate a proportional number of candidate question-answer pairs, explicitly filtering out cases with missing or uncertain information. These candidates are then ranked by the LLMs according to logical consistency, linguistic fluency, semantic relevance, and informational value. Subsequently, multiple rounds of human-assisted review and prompt adjustment are employed to eliminate any hallucinated or fabricated content produced by the LLMs. Finally, a concise set of six diverse and content-rich QA pairs is selected for each video, ensuring comprehensive coverage from broad contextual understanding to detailed actions.
@@ 4 Dataset Composition
Fine-Grained Human Action Recognition Subset: Contains fine-grained, multi-label annotations for spatiotemporal human action detection, comprising $4,759$ manually annotated three-second video clips from authentic microgravity environments. Each clip includes annotations for up to five distinct actions per detected individual. The $50$ action labels are organized into three categories: Object Manipulation ($4,976$ annotations, $37.60\%$), Person Interaction ($4,288$ annotations, $32.34\%$), and Person Movement ($3,987$ annotations, $30.07\%$). Figure 2(d) illustrates the distribution of these categories in both real and simulated video clips, while Figure 2(e) shows the distribution of person counts per clip, highlighting the diversity of social configurations. Overall, the subset contains $13,261$ action annotations, including $9,610$ from real footage, $3,651$ from simulated sources, and $390,000$ bounding box annotations.
Video Caption Subset: Comprises $1,238$ detailed, semantically rich descriptions validated against official aerospace agency documentation for astronaut identities, spacecraft locations, actions, appearances, and interactions. Each caption corresponds uniquely to one video clip. The word cloud visualization in Figure 2(a) provides insights into frequently mentioned terms and concepts, illustrating the thematic and semantic distribution of the captions.
Visual Question Answering (VQA) Subset: Includes $7,428$ structured QA pairs, systematically generated and refined via LLMs to ensure linguistic fluency, semantic relevance, and comprehensive coverage of detailed actions and broader context, with each of the $1,238$ video clips associated with up to six diverse QA pairs. Figure 2(b) and (c) show word clouds for questions and answers separately, revealing prevalent inquiry types and common semantic patterns within the dataset.
Figure 2 presents representative video clips from the dataset, demonstrating typical annotation examples and highlighting the visual and semantic diversity of the dataset. More details regarding the dataset can be found in the appendix.

### v4

@@ Abstract
Despite substantial progress in video understanding, most existing datasets are limited to Earth’s gravitational conditions. However, microgravity alters human motion, interactions, and visual semantics, revealing a critical gap for real-world vision systems. This presents a challenge for domain-robust video understanding in safety-critical space applications. To address this, we introduce MicroG-4M, the first benchmark for spatio-temporal and semantic understanding of human activities in microgravity. Constructed from real-world space missions and cinematic simulations, the dataset includes 4,7594{,}759 clips with 13,26113{,}261 action annotations covering $50$ actions, 1,2381{,}238 context-rich captions, and over 7,0007{,}000 question–answer pairs on astronaut activities and scene understanding. MicroG-4M aims to support three core tasks: fine-grained multi-label action recognition, temporal video captioning, and visual question answering, thereby enabling a comprehensive evaluation of both spatial localization and semantic reasoning in microgravity contexts. We establish baselines using state-of-the-art models. All data, annotations, and code are available at https://github.com/lei-qi-233/MicroG-4M.
@@ 1 Introduction
Yuri Gagarin’s historic flight in 1961 marked the beginning of human space exploration. Since then, significant milestones have been achieved, including crewed lunar landings, the continuous operation of the International Space Station (ISS) for over $25$ years, and the participation of more than $650$ individuals in space missions Moskowitz and Wolf (2025). With numerous planned crewed missions in the near future, the frequency and complexity of human activities in space are expected to increase substantially. In this context, ensuring the safety, enhancing the operational efficiency of space missions, and safeguarding the health and well-being of astronauts are of paramount importance. With the rapid advancement of artificial intelligence, the integration of robotic systems aboard spacecraft is anticipated in the near future. These systems will assist astronauts in routine and mission-critical tasks. Consequently, there is a growing demand for the development of deep learning-based scene understanding and action recognition methods tailored specifically for the unique challenges posed by the microgravity environment.
Human action recognition Wang et al. (2023a); Feichtenhofer (2020); Feichtenhofer et al. (2019); Fu et al. (2019), video captioning Chen et al. (2024), and Visual Question Answering (VQA) Hu et al. (2023) are essential for intelligent human-robot collaboration, particularly in space, where precise perception and understanding of astronauts’ actions and their surrounding context are crucial for ensuring operational safety, efficiency, and autonomous assistance under constrained conditions. This capability is crucial for ensuring mission efficiency, enhancing astronaut safety, and providing autonomous assistance in the confined and complex conditions of space habitats. Numerous human action recognition datasets and video caption datasets have been developed, playing an important role in various research areas within computer vision and in industrial applications Camarena et al. (2023); Pareek and Thakkar (2021); Al-Faris et al. (2020); Le et al. (2022). However, most of the existing datasets for video captioning and action recognition are recorded on Earth without microgravity settings, e.g., Kinetics400 Kay et al. (2017), AVA Gu et al. (2018), FineGym Shao et al. (2020), EgoCross Li et al. (2025), and ActivityNet Captions Chen et al. (2019).
@@
To address this gap, we introduce MicroG-4M, a new video benchmark specifically designed for spatio-temporal and semantic understanding of human activities in microgravity. The name “4M” reflects four characteristics: Multi-source (Real mission footage and physically plausible film), Multimodal (RGB + text annotations), Multi-task (HAR, captioning, VQA), and Microgravity. MicroG-4M comprises 4,7594{,}759 three-second video clips drawn from public YouTube footage of real space missions and carefully selected, realistic space-themed films to augment scenario diversity and coverage. The clips span scenes inside the ISS, the Tiangong Space Station, crewed spacecraft cabins, and extravehicular activities. The corpus contains more than 390,000390{,}000 annotated frames with bounding boxes for each visible individual and 13,000+13{,}000+ action labels covering $50$ distinct actions. In addition, it includes 1,2381{,}238 descriptive clip captions created by human annotators and 7,000+7{,}000+ open-ended question–answer pairs aimed at assessing factual, causal, and counterfactual understanding of scenes in microgravity. Both captions and QA annotations were carefully curated through multiple refinement rounds and data selection stages, ensuring high semantic fidelity, contextual accuracy, and relevance to the microgravity setting. Together, these resources enable multi-label spatio-temporal detection, fine-grained action recognition, caption generation, and VQA within a single dataset. To evaluate existing methods, we build the first comprehensive benchmark for these tasks, dubbed MicroG-Bench. We evaluate representative video encoders (SlowFast Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), and MViT Li et al. (2022)) for human action recognition and leading vision-language models (e.g., InternVideo Wang et al. (2022), Gemini 1.5 Pro Reid et al. (2024), GPT-4o Hurst et al. (2024)) for captioning and VQA tasks. Across all evaluated subtasks, state-of-the-art terrestrial models experience significant performance degradation, underscoring the unique challenges posed by microgravity environments, such as arbitrary orientations, floating objects, and confined spacecraft interiors. These results highlight the necessity of dedicated benchmarks to facilitate the advancement of more robust and generalizable AI systems tailored for space applications. MicroG-Bench, therefore, provides a unified, fair yardstick that will guide the development of more robust and generalizable perception and language systems for astronaut assistance and autonomous mission operations.
We summarize our contributions as follows. We introduce MicroG-4M and MicroG-Bench, establishing the first comprehensive benchmark for microgravity video understanding. Beyond providing 4,7594{,}759 richly-annotated clips across 50 actions, we define a rigorous evaluation protocol with standardized splits for HAR, Captioning, and VQA. Crucially, our in-depth empirical analysis provides the core insight: by benchmarking strong baselines, we quantify a persistent performance gap and uncover characteristic failure modes unique to weightlessness. These findings expose the limitations of terrestrial models, positioning MicroG-4M as the foundational standard for developing future microgravity-aware architectures.
@@ 2 Related Work
Datasets for Action Detection, Video Captioning, and VQA. Progress in video understanding has been largely enabled by the introduction of benchmark datasets across action detection, captioning, and VQA. For action detection, early datasets such as UCF101 Soomro et al. (2012) and HMDB51 Kuehne et al. (2011) provided trimmed classification tasks, later extended by Kinetics Kay et al. (2017) and ActivityNet Caba Heilbron et al. (2015) to large-scale, untrimmed settings with diverse action categories. AVA Gu et al. (2018) further introduced spatio-temporal annotations for atomic actions, enabling fine-grained multi-label detection in realistic scenes. Video captioning evolved from MSVD Chen and Dolan (2011) and MSR-VTT Xu et al. (2016), which paired short clips with multiple sentences, to dense captioning in untrimmed videos via ActivityNet Captions Krishna et al. (2017) and procedural datasets like YouCook2 Zhou et al. (2018). Multilingual benchmarks such as VATEX Wang et al. (2019) expanded the task to cross-lingual settings with high-quality parallel annotations. In VQA, static image datasets like VQA v2.0 Goyal et al. (2017) and CLEVR Johnson et al. (2017) laid the foundation for compositional reasoning, while TGIF-QA Jang et al. (2017), MovieQA Tapaswi et al. (2016), and TVQA Lei et al. (2018) introduced temporal and multimodal reasoning in video. Recent efforts such as NExT-QA Xiao et al. (2021) and CLEVRER Yi et al. (2020) focus on causal inference and counterfactual reasoning. Existing datasets focus on terrestrial settings and lack the unique complexity of space activities. To address this, we introduce the first microgravity benchmark for video captioning and VQA, enabling evaluation in long-horizon, space-relevant scenarios.
@@ 3 Collection Methodology
We aim to construct a dataset comprising authentic footage recorded aboard the International Space Station and other spacecraft, as well as films that realistically depict microgravity conditions. In the following, we introduce our comprehensive pipeline for collecting, assembling, filtering, and annotating the MicroG-4M dataset, which supports multi-label spatio-temporal action detection, video captioning, and video question answering tasks from Internet-sourced videos. An overview of this collection and annotation pipeline is shown in Figure 2.
Figure 2: Overview of the MicroG-4M collection and annotation pipeline. Blue, orange, and purple modules denote automated, human, and MLLM-assisted stages. Raw footage is trimmed into 3-second clips, automatically filtered, and manually verified to retain authentic microgravity scenes. Motion-driven detection produces person-level boxes for HAR, while a real-footage protocol constrains captioning and VQA to genuine microgravity content. AVA-based action labels, along with caption and VQA generation and validation, produce the final annotations for action recognition, captioning, and VQA.
Raw Video Information Collection.
Video sources primarily include genuine microgravity footage from actual spacecraft missions and selected cinematic clips known for their realistic depictions of weightlessness. Authentic spacecraft videos were mainly retrieved from online video platforms, while cinematic clips were manually chosen based on their fidelity to real microgravity conditions. Videos were downloaded with a target resolution of 480480p whenever available, and otherwise at the highest available resolution below 480480p. The final dataset comprises approximately $5,000$ three-second clips, predominantly composed of authentic space station footage.
Dataset Assembly Pipeline.
We established an automated pipeline to preprocess and structure raw videos, enabling consistent and accurate downstream annotation. The pipeline consists of three main stages: video segmentation, filtering, and automated bounding-box annotation.
@@
Filtering. Video clips undergo automatic filtering based on person detection and scene-transition analysis. Specifically, we employed YOLOv11 Khanam and Hussain (2024) for human detection and PySceneDetect Castellano to identify abrupt scene changes, discarding clips with insufficient human-action content or disrupted temporal continuity.
Bounding-box Annotation. Person bounding boxes are automatically annotated using YOLOv11 Khanam and Hussain (2024) detection combined with BoT-SORT tracking Aharon et al. (2022). To enhance annotation accuracy, adaptive strategies were employed by assessing video motion intensity using sparse optical flow methods Jeannin and Divakaran (2001); Ali (2013); Szeliski (2010). Annotation parameters were dynamically adjusted accordingly to optimize computational efficiency and annotation precision. The resulting structured annotations include bounding-box coordinates, unique identities, and detection confidence scores.
@@
Caption and VQA Annotation. For caption and VQA annotation, annotators consulted publicly available, authoritative resources, including official crew rosters and astronaut biography databases from major space agencies (e.g., NASA, CNSA, ESA), historical spacecraft module layout diagrams, equipment specification documents, mission operation reports and task logs, and onboard video transcripts and archival footage. These domain resources are systematically used to resolve ambiguities in astronaut identity, cabin context, and equipment usage, and to enforce consistent naming conventions for modules and objects.
For caption annotation, captions are written by human annotators based on frame-by-frame visual inspection of each 3-second clip at $30$ fps. Beyond the core action and object description, captions explicitly encode astronaut identity, cabin location and background spatial layout, fine-grained hand–body–object interactions, equipment appearance and usage, and posture, orientation, and other distinctive visual traits (e.g., helmet and suit state). All factual content is cross-checked against the domain resources above to ensure semantic accuracy. Multimodal Large Language Models (MLLMs) are applied in a subsequent refinement step to improve grammatical fluency, clarity, lexical diversity, and to simplify overly complex sentences into a standard captioning style; all edited captions are re-validated by annotators. These high-quality captions provide strong semantic grounding to support downstream tasks such as action recognition and context-aware retrieval.
For Video Question Answering (VQA), we adopt a caption-grounded, two-stage generation and filtering pipeline guided by Heilman and Smith (2009). Given each refined caption, candidate question–answer pairs are generated via MLLMs using prompt templates designed to elicit variation across multiple reasoning dimensions, including standard Wh-forms, foreground versus background elements, coarse versus fine-grained actions, identity, location and equipment, and temporal or causal relations. One deliberately unanswerable question per clip is optionally included. A refinement pipeline then filters out candidates that rely on sound, latent intent, or non-visual cues. The remaining QA pairs are ranked by the MLLM based on logical consistency, linguistic fluency, semantic relevance, and information value. Annotators review the top-ranked pairs, remove any hallucinated or fabricated content, revise prompts if necessary, and verify that all retained questions are visually grounded or explicitly marked as “Not mentioned” when unanswerable. To ensure comprehensive coverage from broad contextual understanding to detailed actions while maintaining visual fidelity and a controlled reasoning scope, a final set of six diverse QA pairs is retained for each clip.
@@ 4 Dataset Composition
Fine-Grained Human Action Recognition Subset: Contains fine-grained, multi-label annotations for spatiotemporal human action detection, comprising 4,7594{,}759 manually annotated three-second video clips from authentic microgravity environments. Each clip includes annotations for up to five distinct actions per detected individual. The $50$ action labels are organized into three categories: Object Manipulation (4,9864{,}986 annotations, $37.60\%$), Person Interaction (4,2884{,}288 annotations, $32.34\%$), and Person Movement (3,9873{,}987 annotations, $30.07\%$). Object Manipulation includes actions involving physical objects or equipment, Person Interaction involves social or physical engagement with others, and Person Movement covers individual postural or locomotor motion. These categories are derived from AVA’s Gu et al. (2018) atomic action labels and regrouped for semantic coherence under microgravity conditions, following macro-level groupings similar to those adopted in AVA-based fine-grained action recognition benchmarks Peng et al. (2024). Figure 3(d) illustrates the distribution of these categories in both real and simulated video clips, while Figure 3(e) shows the distribution of person counts per clip, highlighting the diversity of social configurations. Overall, the subset contains 13,26113{,}261 action annotations, including 9,6109{,}610 from real footage, 3,6513{,}651 from simulated sources, and 390,000390{,}000 bounding box annotations. Video Caption Subset: Comprises 1,2381{,}238 detailed, semantically rich descriptions validated against official aerospace agency documentation for astronaut identities, spacecraft locations, actions, appearances, and interactions. Each caption corresponds uniquely to one video clip. The word cloud visualization in Figure 3(a) provides insights into frequently mentioned terms and concepts, illustrating the thematic and semantic distribution of the captions. Visual Question Answering (VQA) Subset: Includes $7,428$ structured QA pairs, systematically generated and refined via MLLMs to ensure linguistic fluency, semantic relevance, and comprehensive coverage of detailed actions and broader context, with each of the $1,238$ video clips associated with up to six diverse QA pairs. Figure 3(b) and (c) show word clouds for questions and answers separately, revealing prevalent inquiry types and common semantic patterns within the dataset. Figure 3 presents representative video clips from the dataset, demonstrating typical annotation examples and highlighting the visual and semantic diversity of the dataset.